40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 11 02:25:11 2012, yuta, Update, LSC, PRMI beam spot motion might be from PRM/BS motion POPDCASDC4.pngFeedback4.pngOLPITYAWandPOPDC4.pngOLPITYAWandPOPDCunlocked.png
    Reply  Wed Jul 11 03:53:41 2012, yuta, Update, LSC, BS 3.3 Hz motion on MI REFLDCASDCMIlocked.png
Message ID: 6955     Entry time: Wed Jul 11 03:53:41 2012     In reply to: 6954
Author: yuta 
Type: Update 
Category: LSC 
Subject: BS 3.3 Hz motion on MI 

It is not as dramatic as PRMI, but I could see BS 3.3 Hz motion at AS and REFL when MI is locked at dark fringe.
Below is uncalibrated spectra of REFLDC and ASDC when
  Red: MI is locked at dark fringe
  Blue: there's no light (PSL shutter closed)

We have to do something to get rid of this.

REFLDCASDCMIlocked.png

ELOG V3.1.3-